# MergeMix: A Unified Augmentation Paradigm for Visual and Multi-Modal Understanding

### v1

@@ B.3 Additional Results of Image Classification
Table 9 shows the full results of 200 epochs and 600 epochs training on the CIFAR100 dataset using different ViT models. Table 10 shows the three fine-grain datasets, i.e., CUB200, FGVC-Aircrafts, and Stanford-Cars. It is easily found that MergeMix can achieve the SOTA on lots of models. Also, the speed of overfitting is a significant improvement over other methods, which means the token merge can gather useful information and reduce some redundant tokens.
B.4 Robustness Experiments of Mixup Augmentations
Calibration of Mixup Augmentations
Table 11 shows the results of calibration on seven mixup augmentations. We evaluated with two public datasets by DeiT-Tiny, ViT-Small, and ViT-Large on the CIFAR100 dataset, by models trained with 200 epochs, DeiT-Tiny and ViT-Small trained with 600 epochs, and DeiT-Small on the ImageNet-1K dataset, trained with 300 epochs.
@@ B.4 Robustness Experiments of Mixup Augmentations
Table 11: The calibration results of ViT-based mixup methods on CIFAR-100 & ImageNet-1K, with training 200 and 300 epochs respectively. tome denotes inference with token merging.
@@
Figure 7: Robustness against image occlusion classification results with different occlusion ratios for different mixup methods based on DeiT-Tiny (left) and DeiT-Small (right) on CIFAR100 and ImageNet-1K datasets.
Results of Occlusion Robustness
The full results of occlusion robustness classification on MergeMix and other mixup methods. Figure 7 show the corve of MergeMix and other mixup method on CIFAR100 dataset and ImageNet-1K dataset. Table 12 and Table 14 show the accuracy results on the CIFAR100 dataset by vanilla and 14 different mixup approaches. Table 13 shows the results of 8 different methods on the ImageNet-1K dataset.
Table 12: The Top-1 accuracy of DeiT-Tiny trained by various Mixup approaches on the CIFAR100 dataset with different occlusion ratios. tome denotes inference with token merging.
@@
Table 13: The Top-1 accuracy of DeiT-Small trained by various Mixup approaches on the ImageNet-1K dataset with different occlusion ratios. tome denotes inference with token merging.
Method DeiT-Samll Trained 300 epochs
0% 10% 20% 30% 40% 50% 60% 70% 80% 90%
Vanilla 75.46 74.03 72.85 71.36 69.91 67.94 64.71 60.42 51.65 34.08
MixUp 78.74 78.17 77.11 76.21 74.31 71.85 68.13 62.40 51.45 33.08
CutMix 80.16 79.23 78.20 76.87 75.21 72.92 69.24 64.63 55.95 39.69
TransMix 80.36 79.47 78.24 77.00 75.40 73.31 70.17 65.78 57.79 42.44
DeiT 80.27 79.03 77.92 76.39 74.65 72.25 68.31 63.53 54.57 38.81
SMMix 79.32 78.56 77.70 76.17 73.76 70.25 64.95 58.94 48.69 34.37
MixPro 79.25 78.83 78.01 77.24 76.02 74.26 71.48 67.44 58.49 39.47
MergeMix 80.70 80.38 79.95 78.97 78.08 76.21 74.08 70.71 65.57 51.98
[…]
Table 14: The Top-1 accuracy of ViT-Small trained by various Mixup approaches on the CIFAR100 dataset with different occlusion ratios. tome denotes inference with token merging.
@@
Appendix C Appendix for MLLMs
C.1 Results of Calibration on MLLMs
Table 15: The calibration results of LLaVA on POPE, ScienceVQAI, GQA & SEEDI.
@@ C.1 Results of Calibration on MLLMs
For exploring the calibration of MLLMs, we chose 4 tasks with a short response under different token reduction settings. We evaluate three scenarios: using full vision tokens, 50% tokens, and 25% tokens, and compare various data augmentation strategies (MixUp, CutMix, ResizeMix, MergeMix) trained with ranking loss (rl). Table 15 shows that with unfreezing the vision encoder, the ECE can be better than freezing. With the vision tokens reduced in training. Overall, token reduction leads to a moderate drop in calibration accuracy, but effective augmentation strategies significantly mitigate this degradation. In the full-token setting, CutMixrl achieves the lowest GQA calibration error (6.09), while ResizeMixrl shows the best SEED result (36.08). When reducing tokens to 50%, CutMixrl and MergeMixrl remain competitive, maintaining strong calibration across tasks despite reduced visual information. Even with 25% tokens, CutMixrl continues to yield relatively balanced performance, indicating that appropriate augmentation enhances robustness under token compression. These results suggest that the calibration of MLLMs is still an open question, especially in some environments that need a reliable answer.
C.2 Results of Different Vision Token Ratios on Inference
@@
In this subsection, we validate the different ratios of vision tokens on the LLaVA benchmark. Table 16, Table 17 and Table 18 show the fully results with full, 75%, 50% and 25% ratios on inference, respectively. Those results give a full comparison of the influence on the vision tokens. Significantly shown in Table 16, MergMix always brings gains in different merge ratios, from +0.83 to +0.06. Other methods, since they are highly random, cause performance instability.
@@
From the results shown in Table 17, when the training stage uses the token merge, it can achieve an average gain of 66.84%, which improves 0.43% over training and inference without token merge training and inference, with an improvement of +1.27% than the average performance of the original LLaVA model.
@@
Table 16: Different token merge ratios of inference comparison results with augmentations. AVG: The average of the nine benchmarks for comprehensive comparison, except for MME.
@@
Table 17: Different token merge ratios of inference comparison results with augmentations. AVG: The average of the nine benchmarks for comprehensive comparison, except for MME.
@@
Table 18: Different token merge ratios of inference comparison results with augmentations. AVG: The average of the nine benchmarks for comprehensive comparison, except for MME.
@@
Appendix D Visualization of MergeMix
In this section, we plot some visualizations of token merge with different merge ratios, mixed samples with different $\lambda$ in Figure 9 and Figure 10, and some case studies of different augmentation degrees in Figure 8.
@@ Appendix D Visualization of MergeMix
Figure 8: The visualization of the visual question answers with different mixing ratios by LLaVA-v1.5-7B model. Note that the blue texts denote the core question and the corresponding correct answers, while the green texts denote the wrong answer to the question. The raw image denotes without any augmentations, and other images denote with different mixing ratios $\lambda$. Ground-truth Answer denotes the raw labels for this case. With the mixing degree improving, the answer comes out more wrong or unrelated to the question, as shown in green color.
@@
Figure 9: The visualization of ToMe with different merge ratios and mixed samples with different mixing ratios.
@@
Figure 10: The visualization of ToMe with different merge ratios and mixed samples with different mixing ratios.

### v2

@@ C.1 Extensive Results
Table A1 shows the full results of 200 epochs and 600 epochs training on the CIFAR100 dataset using different ViT models. Table A2 shows the three fine-grain datasets, i.e., CUB200, FGVC-Aircrafts, and Stanford-Cars. It is easily found that MergeMix can achieve the SOTA on lots of models. Also, the speed of overfitting is a significant improvement over other methods, which means the token merge can gather useful information and reduce some redundant tokens.
@@
Table A3: The calibration results of ViT-based mixup methods on CIFAR-100 & ImageNet-1K, with training 200 and 300 epochs respectively. tome denotes inference with token merging.
@@
Table A4: The Top-1 accuracy of DeiT-Tiny trained by various Mixup approaches on the CIFAR100 dataset with different occlusion ratios. tome denotes inference with token merging.
@@
Table A5: The Top-1 accuracy of ViT-Small trained by various Mixup approaches on the CIFAR100 dataset with different occlusion ratios. tome denotes inference with token merging.
@@
Table A6: The Top-1 accuracy of DeiT-Small trained by various Mixup approaches on the ImageNet-1K dataset with different occlusion ratios. tome denotes inference with token merging.
Method DeiT-Samll Trained 300 epochs
0% 10% 20% 30% 40% 50% 60% 70% 80% 90%
Vanilla 75.46 74.03 72.85 71.36 69.91 67.94 64.71 60.42 51.65 34.08
MixUp 78.74 78.17 77.11 76.21 74.31 71.85 68.13 62.40 51.45 33.08
CutMix 80.16 79.23 78.20 76.87 75.21 72.92 69.24 64.63 55.95 39.69
TransMix 80.36 79.47 78.24 77.00 75.40 73.31 70.17 65.78 57.79 42.44
DeiT 80.27 79.03 77.92 76.39 74.65 72.25 68.31 63.53 54.57 38.81
SMMix 79.32 78.56 77.70 76.17 73.76 70.25 64.95 58.94 48.69 34.37
MixPro 79.25 78.83 78.01 77.24 76.02 74.26 71.48 67.44 58.49 39.47
MergeMix 80.70 80.38 79.95 78.97 78.08 76.21 74.08 70.71 65.57 51.98
[…]
C.2 Robustness Experiments of Mixup Augmentations
Calibration of mixup augmentations
Table A3 shows the results of calibration on seven mixup augmentations. We evaluated with two public datasets by DeiT-Tiny, ViT-Small, and ViT-Large on the CIFAR100 dataset, by models trained with 200 epochs, DeiT-Tiny and ViT-Small trained with 600 epochs, and DeiT-Small on the ImageNet-1K dataset, trained with 300 epochs.
Results of occlusion robustness
The full results of occlusion robustness classification on MergeMix and other mixup methods. Figure A1 shows the curve of MergeMix and other mixup methods on the CIFAR100 dataset and ImageNet-1K dataset. Table A4 and Table A5 show the accuracy results on the CIFAR100 dataset by vanilla and 14 different mixup approaches. Table A6 shows the results of 8 different methods on the ImageNet-1K dataset.
Table A7: Classification results of CAFormer small (CAFormer-S12) with different mxiup augmentations, training 200 epochs with 100 batch size on CIFAR100 dataset.
Model Dataset Epochs Vanila MixUp CutMix DeiT TransMix MergeMix
CAFormer-S12 CIFAR100 200 74.95 81.64 84.69 83.60 83.70 84.30
C.3 Further results on different backbone and data modality.
Table A8: Supervised fine-tuning on HuBERT-Base model on ESC-50 and UbranSound8K datasets.
HuBERT-Base ESC-50 UbranSound8K
Vanilla 75.12±1.07 84.14±0.45
MixUp 75.86±0.83 85.02±0.26
TransMix 76.27±1.14 85.33±0.57
MergeMix 76.51±0.95 85.69±0.42
For further exploring the effectiveness of MergeMix, we applied our approach on a hyper-model MetaFormer (CAFormer) (yu2022metaformer) and two audio datasets for classification by HuBERT-Base (hsu2021hubert). ESC-50 (piczak2015esc) consisted of 50 classes, with 1,200 training samples and 400 validation samples, and a maximum duration of 3 seconds. UrbanSound8k (salamon2014urban8k) is a classification dataset consisting of 10 classes, with a maximum duration of 4 seconds, containing 7,079 training samples and 816 validation samples. Following the USB experimental setup, we fine-tuned the model for 100 epochs using the AdamW optimizer. The base learning rate was set to 1e-4, texttt5e-4, with a batch size of 32 and a weight decay coefficient of 5e-4. Both the compared shuffling methods and our proposed MergeMix can be directly transferred to audio data (treated as one-dimensional sequences). Table A8 below shows our reproduced comparison results on two audio datasets. Compared to the MixUp and TransMix baselines, MergeMix achieves significant performance improvements over multiple shuffling baseline models. For CAFormer-S12, we train for 200 epochs on the CIFAR100 dataset with the same settings as DeiT-Tiny. Table A7 shows that MergeMix obtained a second-best result, since CAFormer only has with Attention module of 2 stages (8 layers).
Appendix D Visual Understanding
D.1 Calibration Results of MLLM
Table A9: The calibration results of LLaVA on POPE, ScienceVQAI, GQA & SEEDI.
@@ D.1 Calibration Results of MLLM
To explore MLLM calibration, we selected 4 tasks with short responses across different token-reduction settings. We evaluate three scenarios: using full vision tokens, 50% tokens, and 25% tokens, and compare various data augmentation strategies (MixUp, CutMix, ResizeMix, MergeMix) trained with ranking loss (denoted as rl). Table A9 shows that with unfreezing the vision encoder, the ECE can be better than freezing. With the vision tokens reduced in training. Overall, token reduction leads to a moderate drop in calibration accuracy, but effective augmentation strategies significantly mitigate this degradation. In the full-token setting, CutMixrl achieves the lowest GQA calibration error (6.09), while ResizeMixrl shows the best SEED result (36.08). When reducing tokens to 50%, CutMixrl and MergeMixrl remain competitive, maintaining strong calibration across tasks despite reduced visual information. Even with 25% tokens, CutMixrl continues to yield relatively balanced performance, indicating that appropriate augmentation enhances robustness under token compression. These results suggest that MLLM calibration remains an open question, especially in environments that require a reliable answer.
D.2 Relationship between Mixing Ratios and Rewards
Figure A2: LLM-as-the-Judge results for different mixing ratios. Higher mixing ratios consistently yield higher reward scores across strong MLLMs. MergeMix aligns closely with these models, showing that the mixing ratio provides a meaningful preference signal.
To understand how the proposed ranking loss on synthetic mixed pairs approximates human preference learning, we conduct a LLM-as-the-judge-eval evaluation in which strong MLLMs score the mixed pairs generated with different mixing ratios $\lambda$. Specifically, we query a diverse set of frontier MLLMs, including Grok-3, Doubao-Seed-1.6 (guo2025seed1), Doubao-Seed-1.6-CoT, Qwen3-VL-Plus-32B (yang2025qwen3), Qwen2.5-VL-72B (bai2025qwen2.5), Gemini-2.5 (team2023gemini), and a human-expert baseline, to assign reward scores to each mixed pair. As shown in Figure A2, we found that the reward consistently increases with larger mixing ratios, and this trend holds across nearly all evaluators. The strong monotonic correlation suggests that $\lambda$ provides a reliable and well-behaved control signal for modeling preference strength. This observation validates our use of $\lambda$ as an interpretable proxy for preference supervision and highlights its potential as a lightweight alternative to explicit human-annotated reward signals.
D.3 Results of Different Vision Token Ratios on Inference
@@
In this subsection, we validate the different ratios of vision tokens on the LLaVA benchmark. Table A10, Table A11 and Table A12 show the fully results with full, 75%, 50% and 25% ratios on inference, respectively. Those results give a full comparison of the influence on the vision tokens. Significantly shown in Table A10, MergMix always brings gains in different merge ratios, from +0.83 to +0.06. Other methods, since they are highly random, cause performance instability.
@@
From the results shown in Table A11, when the training stage uses the token merge, it can achieve an average gain of 66.84%, which improves 0.43% over training and inference without token merge training and inference, with an improvement of +1.27% than the average performance of the original LLaVA model. Figure A3 shows the average accuracy on the LLaVA benchmark for the LLaVA-v1.5-7B model trained and evaluated under different vision token-merging ratios. The results demonstrate that MergeMix maintains strong performance across a wide range of settings, outperforming or matching other baselines.
@@
Table A10: Different token merge ratios of inference comparison results with augmentations. AVG: The average of the nine benchmarks for comprehensive comparison, except for MME.
@@
Table A11: Different token merge ratios of inference comparison results with augmentations. AVG: The average of the nine benchmarks for comprehensive comparison, except for MME.
@@
Table A12: Different token merge ratios of inference comparison results with augmentations. AVG: The average of the nine benchmarks for comprehensive comparison, except for MME.
@@
Figure A3: The plots of the LLaVA-v1.5-7B model under different inference time merge ratios for various methods (Baseline, SFT, MixUp, CutMix, ResizeMix, and MergeMix) demonstrate that MergeMix maintains robust performance across a wide range of configurations.
D.4 Results of Different Ranking Loss
To understand the effectiveness of ranking loss on preference tuning, we conducted an ablation study for $\mathcal{L}^{\text{Mix}}_{\text{SimPo}}$. Table A13 shows that compared with vanilla SimPO, our approach can bring more improvement.
Table A13: Ablation study of different ranking loss on LLaVA-v1.5-7B.
Models SciVQAI TextVQA VizWiz MMBench Avg. Gains
LLaVA-v1.5-7B 66.8 58.2 50.0 64.3 59.82 -
mDOP (wang2024mdpo) 67.53 57.90 50.04 64.60 60.02 +0.20
Re-Align (xing2025re) 68.10 58.55 50.06 64.69 60.35 +0.53
vanilla SimPO 69.86 56.62 49.26 66.24 60.49 +0.67
+ MergeMix 69.86 57.56 47.69 66.58 61.29 +1.47
Appendix E Efficiency
E.1 Results of Different Vision Token Ratios on Inference
To further validate the inference efficiency gains achieved by Token Merge in MergeMix, we conducted experiments on both image classification models and multi-modal large models. As shown in Table A14, increasing the merging ratio $r$ of the ViT-L model significantly reduces FLOPs (from 59.57G to 34.93G) while throughput improves from 122.83 to 201.07 (+63.7%). Moreover, the additional overhead of Token Merge itself is extremely low (only 0.97 ms at $r$ = 0.75), far below the computational cost of pre-layering. This demonstrates that Token Merge can efficiently compress visual tokens while maintaining negligible overhead. For multi-modal inference, Table A15 shows the effectiveness of Token Merge on LLaVA-v1.5-7B. As $r$ increases, the model’s overall throughput improves, while TTFT decreases significantly. The optimal trade-off is achieved at $r$ = 0.5, boosting throughput from 45.96 to 47.28 and reducing TTFT from 86.44 ms to 66.74 ms. When $r$ = 0.75, TTFT further decreases to 62.25 ms. This demonstrates that Token Merge effectively accelerates multi-modal inference without compromising image-text alignment quality. These results collectively show that Token Merge can efficiently reduce the number of visual tokens in MergeMix, thereby lowering FLOPs and inference latency in both visual models and MLLMs. This validates the necessity and advantages of introducing Token Merge into our methodology.
Table A14: Results of throughput, FLOPs (G), overhead of pre-layer and ToMe by ViT-L model with different merged ratios, evaluated on an Nvidia A100 GPU.
Ratios Throughput $\uparrow$ FLOPs (G) Overhead
Layer ToMe
Baseline 122.83 59.57 10.19 -
$r$ = 0.1 130.45 56.29 9.82 3.72
$r$ = 0.25 145.98 49.37 8.23 3.14
$r$ = 0.5 167.54 42.65 5.52 1.79
$r$ = 0.75 201.07 34.93 3.23 0.97
Table A15: Results of throughput and Time-To-First-Token (TTFT) in ms on LLaVA-v1.5-7B with different merged ratios $r$, which is evaluated on a Nvidia A100 GPU.
Ratios Throughput $\uparrow$ TTFT (ms)
LLaVA-V1.5 45.96 86.44
$r$ = 0.1 43.94 89.32
$r$ = 0.25 46.99 83.65
$r$ = 0.5 47.28 66.74
$r$ = 0.75 46.92 62.25
Appendix F Visualization and Case Study
In this section, we provide a visualization of the case study of augmentation samples with corresponding performance reward scores and an extensive visualization of mixing samples and ToMe source maps. Firstly, we provide a case of different degrees of augmentation in Figure A4. Then, we plot some visualizations of token merge with different merge ratios, mixed samples with different $\lambda$ in Figure A5 and Figure A6. For every three rows in Figure A5 and Figure A6, the first and second rows of source maps could directly capture the important regions of the raw images, where a large merge ratio enables better grouping of the similar regions. Based on the source maps, when the mixing ratio $\lambda$ grows from small to large, MergeMix could keep the most distinguishable tokens of the first image while gradually expanding more tokens from less important regions, which enables MergeMix to generate reliable mixing samples with no more computational costs. Moreover, we provide GradCAM (selvaraju2017gradcam) visualization of the top-1 and top-2 classes with the mixed samples of MergeMix on ImageNet-1K, as shown in Figure A7 and Figure A6.
@@ Appendix F Visualization and Case Study
Figure A4: The visualization of the visual question answers with different mixing ratios by LLaVA-v1.5-7B model. Note that the blue texts denote the core question and the corresponding correct answers, while the green texts denote the wrong answer to the question. The raw image denotes without any augmentations, and other images denote with different mixing ratios $\lambda$. Ground-truth Answer denotes the raw labels for this case. With the mixing degree improving, the answer comes out more wrong or unrelated to the question, as shown in green color.
@@
Figure A5: Visualization of mixed samples with source maps of ToMe with different mixing ratios $\lambda$ and various merge ratios on ImageNet-1K.
@@
Figure A6: Visualization of mixed samples with source maps of ToMe with different mixing ratios $\lambda$ and various merge ratios on ImageNet-1K.
Figure A7: Visualization of mixed samples and corresponding GradCAM (selvaraju2017gradcam) of the top-1/2 class with MergeMix on ImageNet-1K.
Figure A8: Visualization of mixed samples and corresponding GradCAM (selvaraju2017gradcam) of the top-1/2 class with MergeMix on ImageNet-1K.